40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 26 17:07:41 2019, gautam, Update, LSC, POX / POY calibration POX_POY_sensorNoise.pdfALSnoise_20191125.pdf
    Reply  Tue Nov 26 18:16:08 2019, rana, Update, LSC, POX / POY calibration 
       Reply  Wed Nov 27 12:16:52 2019, gautam, Update, LSC, ITMX and ITMY OSEMs with low and high circulating power ITMXshadowSensors.pdfITMYshadowSensors.pdf
          Reply  Wed Nov 27 13:14:02 2019, rana, Update, LSC, ITMX and ITMY OSEMs with low and high circulating power 
             Reply  Wed Nov 27 16:10:29 2019, gautam, Update, LSC, AOM reconnected 
Message ID: 15052     Entry time: Wed Nov 27 13:14:02 2019     In reply to: 15051     Reply to this: 15053
Author: rana 
Type: Update 
Category: LSC 
Subject: ITMX and ITMY OSEMs with low and high circulating power 

if the RP don't fit

u must acquit

sweep the laser amplitude

to divine the couplin w certitude

ELOG V3.1.3-